OpenStudy (anonymous):

Please I need help on H4P1 and H4P2 equations, I have been having problem solving for vo.

OpenStudy (anonymous):

Given: Vin, vin, RL,Rin) for the first part do KVL. so you will get Vo=Vin(RL*Rin)/(RL+Rin) for zener diode we look at the graph and notice there are 3 parts. And we know that it can not be operating in the middle. also it can't be on the right since this is a small signal region so it must be on the left. (Note that the graph is pointing down and the diode is pointing up so the graph is upside down) so to solve for VO we look at the graph and see that it is equal to 5 to solve for vo we need to use the equation vo=vin/Rin *(RD*RL)/(RD+RL) (Note: RD is 1ohm from the slope on the graph) and finally to solve for RL we use the equation Vo=Vin*RL/(RL+RIN) Hope this helps

OpenStudy (anonymous):

check out http://openstudy.com/study#/updates/4f823036e4b0505bf083a27f for H4P2

OpenStudy (anonymous):

Also if you get stuck you can check out https://6002x.mitx.mit.edu/discussion/question/39177/hints-on-hw4